OpenStudy (yamyam70):

Jupiter rotates every 9.9 hours , diameter of 88, 846 miles. if you're standing on its equator, how fast are you traveling? Can anyone help me getting the correct formula for this?

ganeshie8 (ganeshie8):

find angular velocity \(\omega \) after that multiply it wid radius \(r\)

ganeshie8 (ganeshie8):

\(v = r \omega \)

OpenStudy (yamyam70):

Thanks again @ganeshie8 :))

ganeshie8 (ganeshie8):

np :)

OpenStudy (yamyam70):

@ganeshie8 the formula for the angular velocity is given by, \[\omega = \frac{ \Theta }{ t } , \] CAn you help me with getting theta ? :)

ganeshie8 (ganeshie8):

here its taking 9.9 hours for one rotation right ?

OpenStudy (yamyam70):

Yes

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

so, \(\theta = 2 \pi\) \(t = 9.9\) hours

OpenStudy (yamyam70):

Thank you very much :) , again :))

ganeshie8 (ganeshie8):

you wlcme :)